OpenStudy (anonymous):

(4+3 square root 3)(9-2 square root 3)

OpenStudy (radar):

\[(4+3\sqrt{3})(9-2\sqrt{3})\]

OpenStudy (radar):

\[36+19\sqrt{3}-18=18+19\sqrt{3}\]